# Deep Learning solutions to singular problems: from special functions to spherical accretion

### v1

@@ Abstract
Singular regular points often arise in differential equations describing physical phenomena such as fluid dynamics, electromagnetism, and gravitation. Traditional numerical techniques often fail or become unstable near these points, requiring the use of semi-analytical tools, such as series expansions and perturbative methods, in combination with numerical algorithms; or to invoke more sophisticated methods. In this work, we take an alternative route and leverage the power of machine learning to exploit Physics Informed Neural Networks (PINNs) as a modern approach to solving differential equations with singular points. PINNs utilize deep learning architectures to approximate solutions by embedding the differential equations into the loss function of the neural network. We discuss the advantages of PINNs in handling singularities, particularly their ability to bypass traditional grid-based methods and provide smooth approximations across irregular regions. Techniques for enhancing the accuracy of PINNs near singular points, such as adaptive loss weighting, are used in order to achieve high efficiency in the training of the network. We exemplify our results by studying four differential equations of interest in mathematics and gravitation – the Legendre equation, the hypergeometric equation, the solution for black hole space-times in theories of Lorentz violating gravity, and the spherical accretion of a perfect fluid in a Schwarzschild geometry.
@@ IV.1 Legendre Differential Equation
We performed several experiments to test the robustness of our approach. The hyperparameters used are summarized in table 1. Benchmark results for the first few Legendre polynomials are shown in figure 2. We observe that at a fixed size of the network and training epochs, the result of the PINN is worse as the solution becomes more complex. We can test the origin of this dependence by increasing the number of hidden layers and neurons – cf. figure 3. We observe that while a large number of layers yields a substantial reduction of the loss value, there is not a clear gain by increasing the number of neurons. All experiments were performed with $M=50$ collocation points. Fluctuations in the value of the loss correspond to the resampling of the collocation points, which nonetheless allows for an improvement of the training by decreasing its trend value.
@@ IV.2 Hypergeometric Equation
(a) Some hypergeometric functions, obtained using PINNs.
@@
Figure 4: Benchmark results for the hypergeometric equation (17). All solutions were obtained with a NN with a single $H=2$ hidden layer and $O=64$ neurons. The vertical dashed line in magenta denotes the position of the singular point $x=1$.
The set of hyperparameters chosen for these experiments is shown in table 2, while benchmark results for a bunch of solutions can be found in figure 4. Note that all of them are perfectly smooth through the singular point at $x=1$, but convergence worsens as we move apart from it. We believe that this is simply a border effect which can be alleviated by enlarging the integration regime and the number of collocation points, as well as allowing for larger training times. However, such fine study is beyond the scope of this work. Note that here, unlike in the case of the Legendre equation (14), we find it more efficient to work with shallow networks but a larger number of neurons – see figure 5.
@@
Figure 5: Loss function in terms of the number of epochs for solving the hypergeometric equation for $(a,b,c)=(-2,2,2)$ with different number of layers and neurons in the NN.
@@ IV.3 Black holes in Lorentz violating gravity
Figure 6: Results for the Lorentz violating scalar graviton $A$, for different values of the spin-0 propagation speed squared $s_{0}^{2}$, obtained using PINNs, compared to traditional methods. All PINN solutions were obtained with a NN with $H=6$ hidden layers and $O=32$ neurons in each layer. For reference, these solutions are found for $M=0.5$, for which, the spin-0 horizon lie at $x_{s}=1.11$ for $s_{0}^{2}=3$, $x_{s}=1.08$ for $s_{0}^{2}=2.1$, and $x_{s}=1.05$ for $s_{0}^{2}=1.5$.
@@ IV.4 Spherical accretion
Evolution of the loss function for different number of hidden layers $H$, keeping $O=256$ neurons in each layer.
(a) Evolution of the loss function for different number of neurons, in a model with $H=3$ hidden layers.
@@ V Conclusions
Throughout this work, we have investigated whether PINNs can be regarded as a efficient tool to solve differential equations with singular points, which are common in many physics applications. We have done so by exploring four examples of problems with increasing complexity – the Legendre equation, the hypergeometric equation, the solution to black hole configurations in Lorentz violating gravity, and the spherical accretion of a perfect fluid on a Schwarzschild background.
In all cases, we have found that PINNs are able to converge to the right solutions efficiently. Such success is possible by implementing several techniques that increase the efficiency of training – adaptive weights, schedule control of the learning rate through cosine annealing, and specific methods to select the collocation points in the integration regime, such as randomization and domain decomposition. Altogether, these allow us to reach small values of the loss function – for instance, ${\cal L}\sim 10^{-11}$ in the case of spherical accretion – in a matter of seconds when running on a NVIDIA GeForce 4070 Super GPU.
Our results imply that PINNs can be valuable tools to attack numerical problems involving differential equations, either by themselves or in combination with standard methods based on finite differences, which are currently in a much more mature state. One could envision PINNs being used to transverse through singular points, while a traditional algorithm is used to continue evolving the system in the regular region. Alternatively, one could produce data-points by means of standard algorithms and feed them into a PINN as an unsupervised task in order to increase efficiency of the latter. Many more options might be possible, and we expect PINNs to become standard tools in numerical analysis in the coming future.

### v2

# Deep Learning solutions to singular ordinary differential equations: from special functions to spherical accretion

@@ Abstract
Singular regular points often arise in differential equations describing physical phenomena such as fluid dynamics, electromagnetism, and gravitation. Traditional numerical techniques often fail or become unstable near these points, requiring the use of semi-analytical tools, such as series expansions and perturbative methods, in combination with numerical algorithms; or to invoke more sophisticated methods. In this work, we take an alternative route and leverage the power of machine learning to exploit Physics Informed Neural Networks (PINNs) as a modern approach to solving ordinary differential equations with singular points. PINNs utilize deep learning architectures to approximate solutions by embedding the differential equations into the loss function of the neural network. We discuss the advantages of PINNs in handling singularities, particularly their ability to bypass traditional grid-based methods and provide smooth approximations across irregular regions. Techniques for enhancing the accuracy of PINNs near singular points, such as adaptive loss weighting, are used in order to achieve high efficiency in the training of the network. We exemplify our results by studying four differential equations of interest in mathematics and gravitation – the Legendre equation, the hypergeometric equation, the solution for black hole space-times in theories of Lorentz violating gravity, and the spherical accretion of a perfect fluid in a Schwarzschild geometry.
@@ IV.1 Legendre Differential Equation
We performed several experiments to test the robustness of our approach. The hyperparameters used are summarized in table 1. Benchmark results for the first few Legendre polynomials are shown in figure 2. We observe that at a fixed size of the network and training epochs, the result of the PINN is worse as the solution becomes more complex. We can test the origin of this dependence by increasing the number of hidden layers and neurons – cf. figure 3. We observe that while a large number of layers yields a substantial reduction of the loss value, there is not a clear gain by increasing the number of neurons beyond $\sim 32$ neurons per layer, where it reaches saturation. All experiments were performed with $M=50$ collocation points; however, we observe that such saturation is only reached at a higher number of neurons per layer if the number of collocation points is increased. Fluctuations in the value of the loss correspond to the resampling of the collocation points, which nonetheless allows for an improvement of the training by decreasing its trend value.
@@ IV.2 Hypergeometric Equation
(a) Some hypergeometric functions, obtained using PINNs contrasted to the analytical solutions.
@@
Figure 4: Benchmark results for the hypergeometric equation (17). All solutions were obtained with a NN with a single $H=2$ hidden layer and $O=64$ neurons. The vertical dashed lines in magenta denote the positions of the singular points $x=0$ and $x=1$.
The set of hyperparameters chosen for these experiments is shown in table 2, while benchmark results for a bunch of solutions can be found in figure 4. Note that all of them are perfectly smooth through the singular points at $x=0$ and $x=1$. The error of the PINNs solutions tends to grow past the $x=1$ point, we noted that to keep this error under control it was necessary to train the network by employing domain decomposition. If the total spatial domain is given by $\mathcal{D}=\{x_{b},x_{f}\}$, we first train the network in a domain $\mathcal{D}_{i}=\{x_{b},x_{i}\}$, where $x_{i}<x_{f}$, until the loss is below a desired threshold $\mathcal{T}_{\mathcal{L}}$. Subsequently, the training continues in a domain $\mathcal{D}_{i+1}=\{x_{b},x_{i+1}\}$, where $x_{i}<x_{i+1}<x_{f}$. This procedure can be iterated until reaching the desired spatial domain $\mathcal{D}$. Note that here, unlike in the case of the Legendre equation (14), we find it more efficient to work with shallow networks but a larger number of neurons – see figure 5.
@@
Figure 5: Loss function in terms of the number of epochs for solving the hypergeometric equation for $(a,b,c)=(-2,2,2)$ in the domain $x=\{x=0,1.2\}$ with different number of layers and neurons in the NN.
@@ IV.3 Black holes in Lorentz violating gravity
Figure 6: Results for the Lorentz violating scalar graviton $A$, for different values of the spin-0 propagation speed squared $s_{0}^{2}$, obtained using PINNs, compared to traditional methods. All PINN solutions were obtained with a NN with $H=6$ hidden layers and $O=32$ neurons in each layer. For reference, these solutions are found for $M=0.5$, for which, the spin-0 horizon lies at $x_{s}=1.11$ for $s_{0}^{2}=3$, $x_{s}=1.08$ for $s_{0}^{2}=2.1$, and $x_{s}=1.05$ for $s_{0}^{2}=1.5$.
@@ IV.4 Spherical accretion
(a) Evolution of the loss function for different number of hidden layers $H$, keeping $O=256$ neurons in each layer.
(b) Evolution of the loss function for different number of neurons, in a model with $H=3$ hidden layers.
@@ V Conclusions
Throughout this work, we have investigated whether PINNs can be regarded as an efficient tool to solve ordinary differential equations with singular points, which are common in many physics applications. We have done so by exploring four examples of problems with increasing complexity – the Legendre equation, the hypergeometric equation, the solution to black hole configurations in Lorentz violating gravity, and the spherical accretion of a perfect fluid on a Schwarzschild background. In principle, the same PINN-based approach can be generalized to the more involved case of partial differential equations with singular points. [^4] A trivial example is the wave equation in spherical coordinates, where the center is a regular singular point.
In all cases, we have found that PINNs are able to converge to the right solutions efficiently. Such success is possible by implementing several techniques that increase the efficiency of training – adaptive weights, schedule control of the learning rate through cosine annealing, and specific methods to select the collocation points in the integration regime, such as randomization and domain decomposition. Altogether, in the examples explored, these allowed us to reach small values of the loss function in a matter of seconds when running on a NVIDIA GeForce 4070 Super GPU  [^5] The NVIDIA GeForce 4070 Super GPU is far from the raw power found in modern workstation GPUs (for instance, Nvidia H100 series nodes, which are the standard high-efficiency GPUs in modern supercomputers, triple the theoretical performance and VRAM memory of the 4070 Super, while also containing a larger number of more efficient tensor cores).. Moreover, the training time does not change significantly when the hyperparameters of the networks are varied within reasonable values.
Our results imply that PINNs can be valuable tools to attack numerical problems involving differential equations, either by themselves or in combination with standard methods, which are currently in a much more mature state. One could envision PINNs being used to transverse through singular points, while standard integration methods can be employed to seamlessly extend the solution in the regular region in a more cost-effective manner. Alternatively, one could produce data-points by means of standard algorithms and feed them into a PINN as an unsupervised task in order to increase efficiency of the latter. Many more options might be possible, and we expect PINNs to become standard tools in numerical analysis in the coming future.